OpenStudy (anonymous):

How do I integrate (x^2-6)/(x^2-9)?

OpenStudy (blurbendy):

Hm, if substitution doesn't work then try integration by parts. (x^2 - 6)(x^2-9)^-1

OpenStudy (anonymous):

what if there is a partial fraction decomposition of this quotient?

OpenStudy (anonymous):

divide, then use partial fractions

OpenStudy (anonymous):

actually there is an easy gimmick for division write the numerator as \[x^2-9+3\]

OpenStudy (anonymous):

then you get \[\int 1dx-3\int\frac{dx}{x^2-9}\]

OpenStudy (blurbendy):

@satellite73 nice insight

OpenStudy (anonymous):

oops that was wrong \[\int 1dx+3\int\frac{dx}{x^2-9}\]

OpenStudy (anonymous):

@blurbendy thanks second part use partial fractions

OpenStudy (anonymous):

ok thanks for your help. I guess I thought there was a more conventional approach.